When evaluating the question f(4)= -x^2 will it become f(4)=-(4)^2 = -16 OR f(4)=-(4)^2 = 16 ?
Answer by stanbon(75887)

When evaluating the question f(4)= -x^2 will it become f(4)=-(4)^2 = -16 OR f(4)=-(4)^2 = 16 ?
----
If f(x) = -x^2
then f(4) = -(4^2) = -16
===========================
Cheers,
Stan H.
===========